# The MF property for amalgamated free products

### 13 Mar 2026 v1

@@ Abstract
It is proved that for any MF C*-algebra $A$ and its C*-subalgebra $C$, $A\ast_{C}A$ is MF. For general amalgamated free products, $A\ast_{C}B$, a necessary and sufficient condition for being MF is given. It is shown that the following groups – amalgamated free products of amenable groups, semidirect products of amenable groups by free groups, and $\mathbb{Z}^{2}\rtimes SL_{2}(\mathbb{Z})$ – all have MF full group C*-algebra. It is shown that the class of MF C*-algebras is closed under maximal tensor products with $C^{*}(\mathbb{F}_{n})$. We also show that for any Hilbert-Schmidt stable C*-algebra $A$ and any its C*-subalgebra $C$, all hyperlinear traces on $A\ast_{C}A$ are MF.
@@ 1 Introduction
Approximation properties, such as hyperlinearity, quasidiagonality and the MF property, are of central importance in the C*-theory. Similar approximation properties for traces were introduced in [5], and turned out to play an important role in classification of nuclear simple C*-algebras ([30], [24]). It is an open problem whether all hyperlinear traces are MF.
Examples of C*-algebras with all hyperlinear traces being MF include all nuclear simple C*-algebras satisfying UCT ([30]), many crossed product C*-algebras ([22]) and all cone C*-algebras ([26]). Here we find more examples:
Theorem (Cor. 15) Let $A$ be a separable C*-algebra and let $C$ be any C*-subalgebra of $A$. If $A$ is Hilbert-Schmidt stable, then every hyperlinear trace on $A\ast_{C}A$ is MF.
As an easy consequence of the theorem above we obtain that all hyperlinear traces on $C^{*}(\mathbb{F}_{2}\times\mathbb{F}_{2})$ are MF.
@@ 3 Amalgamated free products $A\ast_{C}A$
3.1 MF traces
A trace $\tau$ on a C*-algebra $A$ is hyperlinear if for any $\epsilon>0$ and a finite $G\subset A$ there is $n\in\mathbb{N}$ and a map $\phi:A\to M_{n}$ such that
$\|\phi(ab)-\phi(a)\phi(b)\|_{2}<\epsilon\;\;\text{and}\;\;|\tau(a)-tr\phi(a)|<\epsilon,$
for any $a\in G$.
(Equivalently, one can say that $\tau$ is hyperlinear if $\tau=tr\circ f$, for some $\ast$-homomorphism $f:A\to\prod M_{n}/\oplus_{2}M_{n}$, where $\oplus_{2}M_{k_{n}}$ is the ideal of all sequences that converge to zero in the 2-norm.)
A trace $\tau$ is MF if for any $\epsilon>0$ and a finite $G\subset A$ there is $n\in\mathbb{N}$ and a map $\phi:A\to M_{n}$ such that
$\|\phi(ab)-\phi(a)\phi(b)\|<\epsilon\;\;\text{and}\;\;|\tau(a)-tr\phi(a)|<\epsilon,$
for any $a\in G$.
The property of a trace to be MF is another property that can be formulated in terms of asymptotic lifting [27].
Proposition 13.
([27, Prop. 22]) Let $\tau$ be a trace on a C*-algebra $A$. TFAE:
  • (i)
    $\tau$ is an MF trace,
  • (ii)
    there exists a $\ast$-homomorphism $f:A\to\prod M_{m_{n}}/\oplus_{2,\omega}M_{m_{n}}$ such that $\tau=trf$ and $f$ lifts to a discrete asymptotic homomorphism from $A$ to $\prod M_{n}$,
In particular Proposition 13 implies that if all $\ast$-homomorphisms from $A$ to $\prod M_{m_{n}}/\oplus_{2,\omega}M_{m_{n}}$ lift to discrete asymptotic homomorphisms to $\prod M_{n}$, then all hyperlinear traces on $A$ are MF.
The same arguments as in the proof of Theorem 10 apply to obtain the following result.
Theorem 14.
Let $A$ be a separable C*-algebra and let $C$ be any C*-subalgebra of $A$. If every $\ast$-homomorphism from $A$ to $\prod M_{m_{n}}/\oplus_{2,\omega}M_{m_{n}}$ lifts to a discrete asymptotic homomorphism to $\prod M_{n}$, then the same holds for $A\ast_{C}A$. In particular, every hyperlinear trace on $A\ast_{C}A$ is MF.
A C*-algebra is called Hilbert-Schmidt stable (HS-stable) if every $\ast$-homomorphism from $A$ to $\prod M_{m_{n}}/\oplus_{2,\omega}M_{m_{n}}$ lifts to a $\ast$-homomorphism to $\prod M_{n}$.
Corollary 15.
Let $A$ be a separable C*-algebra and let $C$ be any C*-subalgebra of $A$. If $A$ is Hilbert-Schmidt stable, then every hyperlinear trace on $A\ast_{C}A$ is MF.
Example 16.
$\mathbb{F}_{2}\times\mathbb{F}_{2}$ can be written as amalgamated free product
$\mathbb{F}_{2}\times\mathbb{F}_{2}\cong(\mathbb{F}_{2}\times\mathbb{Z})\ast_{\mathbb{F}_{2}}(\mathbb{F}_{2}\times Z).$
By [12, Th. 2.7] (also [15]) $C^{*}(\mathbb{F}_{2}\times\mathbb{Z})$ is Hilbert-Schmidt stable. Therefore every hyperlinear trace on $C^{*}(\mathbb{F}_{2}\times\mathbb{F}_{2})$ is MF.
@@ 5 Central HNN-extensions and maximal tensor product with $C^{*}(F_{n})$
Similar to the proof of Th. 28. Let $\pi:\langle A,t\;|\;t^{-1}Bt=B,\;id\rangle\to B(H)$ be an embedding. We define a representation $\pi^{\prime}:\langle A,t\;|\;t^{-1}Bt=B,\;id\rangle\to B(H)$ by

### 22 Mar 2026 v2

@@ Abstract
It is proved that for any MF C*-algebra $A$ and its C*-subalgebra $C$, $A\ast_{C}A$ is MF. For general amalgamated free products, $A\ast_{C}B$, a necessary and sufficient condition for being MF is given. It is shown that the following groups – amalgamated free products of amenable groups, semidirect products of amenable groups by free groups, and $\mathbb{Z}^{2}\rtimes SL_{2}(\mathbb{Z})$ – all have MF full group C*-algebra. It is shown that the class of MF C*-algebras is closed under maximal tensor products with $C^{*}(\mathbb{F}_{n})$.
@@ 5 Central HNN-extensions and maximal tensor product with $C^{*}(F_{n})$
Similar to the proof of Th. 24. Let $\pi:\langle A,t\;|\;t^{-1}Bt=B,\;id\rangle\to B(H)$ be an embedding. We define a representation $\pi^{\prime}:\langle A,t\;|\;t^{-1}Bt=B,\;id\rangle\to B(H)$ by